OpenStudy (anonymous):

Write the slope-intercept equation of the line perpendicular to y = 5/2 x - 2, which passes through the point (0, 2).

OpenStudy (zepp):

What would be the slope of a function that's perpendicular to y = 5/2x -2?

OpenStudy (anonymous):

5/2

OpenStudy (anonymous):

?

OpenStudy (zepp):

A slope of a function that's perpendicular to y = 5/2x - 2 would be -2/5 (Change sign, flip the fraction)

OpenStudy (zepp):

Now you have \[y=\frac{-2}{5}x\]

OpenStudy (anonymous):

oh, ok...

OpenStudy (zepp):

To find the y intercept, plug the point into this \[y = \frac{-2}{5}x+b\] Plug the point, (0, 2) \[2 = \frac{-2}{5}0+b\]\[2 = b\]

OpenStudy (anonymous):

how do i do this one? Write the equation for the horizontal line passing through (4, -6).

OpenStudy (zepp):

Well, you have the slope, now what you need to find is the y-intercept so you can satisfy the condition.

OpenStudy (zepp):

Now we know that the y-intercept of this perperndicular function is 2 We put that back, \[y = \frac{-2}{5}x + 2\]